A car is travelling on a straight highway. At 10 o'clock it passes a truck travelling in the same direction. The truck continues on the highway travelling at 50 km/h while the car travelling at a constant speed which is 30% faster than the truck. How far apart are the cars and the truck at 2 o'clock?
Answer by mananth(16946)

From 10 .00 to 2.00 it is 4 hours
time travelled by both is same
speed of truck = 50 km/h
speed of car = 65 km/h ( 30% faster)
distance travelled by truck = 4*50 = 200 km
Distance travelled by car = 4*65 = 260 km
So you have the answer